OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 2x + 10y = 58 -14x + 8y = 140

OpenStudy (anonymous):

2x + 10y = 58 (multiply by 7) -14x + 8y = 140 14x + 70y = 406 -14x + 8y = 140 (x cancels out) 78y = 546 y = 7 plug 7 into any of the original equations to give you the value of x, and your done